OpenStudy (anonymous):

HELP PLEASE! I'll be your fan Part 1. Create two radical equations, one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model. a√x+b+c=d Use a constant in place of each variable a, b, c, and d. You can use positive and negative constants in your equation. Part 2. Show your work in solving the equation. Include the work to check your solution and show that your solution is extraneous. Part 3. Explain why the first equation has an extraneous solution and the second does not.

OpenStudy (anonymous):

lol I suck at everything XD

OpenStudy (anonymous):

@pgpilot326

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

@ehuman @SolomonZelman

OpenStudy (anonymous):

thanks :)

OpenStudy (solomonzelman):

Ok, for part a. \[a√x+b+c=d\] has solution: \[2√x+5+7=4\] has no extraneous solution. \[2√x+5+7=11\] (b/c 2√x=-1, √x=-.5 x=.25. if you plug it in, you will get a decimal for 2√x)

OpenStudy (anonymous):

this is the answer to part 1 ? what about the a b c and d variables and constants or whatever

OpenStudy (solomonzelman):

Yes yes, I used 2 5 7 and 4, as the constants that I plugged in for a b c and d.

OpenStudy (anonymous):

thats it for part 1 ? thats exactly what i should submit ?

OpenStudy (solomonzelman):

At frst you should understand, you should be able t do part 2 and 3, after you have art 1.

OpenStudy (anonymous):

@SolomonZelman lol yea i wish i did understand

OpenStudy (anonymous):

but ill try to do something, can you help me with another question ?

OpenStudy (solomonzelman):

Will see how it goes, I am ding some other staff too.

OpenStudy (anonymous):

ok, ill post it tag you and you get to it when you can or private message me, thanks !

OpenStudy (anonymous):

lol I know this was a long time ago, but was this answer right??

OpenStudy (anonymous):

@yeah_i_suck_at_math @SolomonZelman